OpenStudy (anonymous):

Solve and check. r\6 = 5 2\3 A.34 b.30 2\3 c.20 d. 17\18

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

First, get r alone. To do that, multiply 6 on both sides.

OpenStudy (anonymous):

I don't understand :| I need to get this done fast i have around 30 mins

OpenStudy (anonymous):

@Serenity74

OpenStudy (serenity74):

I can't give you the answer dude. Make the mixed fraction into a fraction and then multiply that fraction by 1/6 and then you get the answer for r.

OpenStudy (serenity74):

Mixed fraction into an improper fraction*

OpenStudy (serenity74):

I know the answer already.

OpenStudy (anonymous):

34?

OpenStudy (serenity74):

No. You are supposed to get a fraction as your answer.

OpenStudy (serenity74):

5 & 2/3 = 17/3 So in order to isolate the r you multiply 1/6 on both sides so you get r = 1/6 * (17/3) so what do you get now?